The Balmer series for the hydrogen atom corresponds to electronic transitions that terminate in the state with quantum number n=2. Consider the photon of longest wavelength corresponding to a transition. Determine (a) its energy and (b) its wavelength.
1
Expert's answer
2020-10-16T10:59:23-0400

Since we need to consider the photon with the longest wavelength that ends the electron transition in the state with the quantum number "n=2" , this will be transition "3-2" since the photon energy formula "E=h\times \frac{c}{\lambda}" , (where "h" is the Planck constant, "c" is the speed of light, and "\lambda" is the wavelength) shows that the photon energy is inversely proportional to the wavelength, and since a larger transition corresponds to a large absorption/release of energy, we choose the smallest transition (transition "1-2>" transitions "2-3" ). By the formula "\frac{1}{\lambda}=R\times(\frac{1}{m^2}-\frac{1}{n^2})" (where "R" is the Rydberg constant, and "n" and "m" are the quantum numbers of the photon state) "\frac{1}{\lambda}=3.29\times10^{15}\times(\frac{1}{2^2}-\frac{1}{3^2})" "=\frac{5\times 3.29\times10^{15}}{36}" therefore "\lambda=\frac{36}{5\times 3.29\times10^{15}}=2.19\times 10^{-15}" m.

"E=6.63\times 10^{-34}\times \frac{3\times 10^8}{2.19\times 10^{-15} }=9.08\times 10^{-11}" J.